OpenStudy (anonymous):

Solve the following system of equations: 2x - y + z = -3 2x + 2y + 3z = 2 3x - 3y - z = -4

OpenStudy (anonymous):

will reward medal and follow

OpenStudy (mrnood):

take this equation: 2x - y + z = -3 add y to each side and add 3 to each side This will give you an equation for y (in terns of x and z) do that first... post result here

OpenStudy (anonymous):

2x+x=-3+y

OpenStudy (anonymous):

@MrNood